WACC Information extracted from XYZ firm are as follow: Common Shares issued: 1 million shares Share...

WACC

Information extracted from XYZ firm are as follow:

Common Shares issued: 1 million shares

Share Price: $5 per share

Bonds issued: 10,000 bonds

Face Value of Bond: $1,000

Price of Bond: $1,044.52

Coupon rate of bond: 5%

Maturity of Bond: 5 years

Information extracted from Factset are as follow:

XYZ’s beta: 1.2 times of market beta

10-year Treasury Bond yield: 3%

5-year AA bond yield: 3.5%

5-year A bonds have a spread of 0.5% above AA bond yields

Expected Return of the stock market: 9%

Tax Rate: 20%

Calculate XYZ’s WACC. (show workings)
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=5*1000000
=5000000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*10000*1.04452
=10445200
MV of firm = MV of Equity + MV of Bond
=5000000+10445200
=15445200
Weight of equity = MV of Equity/MV of firm
Weight of equity = 5000000/15445200
W(E)=0.3237
Weight of debt = MV of Bond/MV of firm
Weight of debt = 10445200/15445200
W(D)=0.6763
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 3 + 1.2 * (9 - 3)
Cost of equity% = 10.2
Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =5
1044.52 =∑ [(5*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^5
                   k=1
YTM = 3.9999611813
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 3.9999611813*(1-0.2)
= 3.19996894504
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=3.2*0.6763+10.2*0.3237
WACC =5.47%